OpenStudy (anonymous):

can x^8-1 can be factored?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

yes.. \[x^8-1=(x^4+1)(x^4-1)=(x^4+1)(x^2+1)(x^2-1)=(x^4+1)(x^2+1)(x+1)(x-1)\] =(x^4+1)(x^2+1)(x+1)(x-1)